OpenStudy (anonymous):

Prove that the given equations are identities: a. (1+tan^2x)(1+cos2x)=2 b. (cosxcscx)/(tanx+cotx)=sin^2x+cos2x Please help with one or both! I've gotten both part way but am messing up somewhere...

mathslover (mathslover):

\(\cos 2x = \cos^2 x - \sin^2 x\) \(1+ \tan^2 x = \sec ^2 x\) Use the above formulas to prove i)

mathslover (mathslover):

Here is how to do it : \(\mathsf{(1+\tan^2 x)(1+\cos 2x) \\ (\sec^2 x)( 1+ \cos^2 x - \sin^2 x) (\sec^2 x)(1 - \sin^2 x + \cos^2 x) \\ (\sec^2 x)(\cos^2 x + \cos^2 x) \\ (\sec^2 x)(2\cos^2 x)\\ }\) Write : \(\sec^2 x = \cfrac{1}{\cos^2 x}\)

mathslover (mathslover):

prove it then.

mathslover (mathslover):

For ii) \(\mathsf{(\cos x \csc x)/(\tan x+\cot x)=\sin^2 x+\cos 2x}\) Hint : Write csc x = 1/sinx tanx = sinx/cos x cot x = cos x / sin x

OpenStudy (anonymous):

And for the second, I had originally gotten cotx/(tanx+cotx)=cos^2x from using the double angle formula and combining the numerator, but I wasn't sure where to go from there

mathslover (mathslover):

\(\cfrac{\cfrac{1}{\tan x} }{\tan x + \cfrac{1}{\tan x} } = \cos^2x\)

mathslover (mathslover):

\(\cfrac{\cfrac{1}{\tan x} }{\cfrac{\tan^2 x + 1}{\tan x}} = \cos^2 x\) \(\cfrac{1}{\sec^2 x} = \cos^2 x \)

mathslover (mathslover):

This is what I did : \(\cfrac{\cfrac{1}{\tan x} . \tan x }{\tan^2 x + 1} \) \(\cfrac{1}{\tan^2 x + 1}\) \(\tan^2 x + 1 = \sec^2 x\) -- identitty \(\cfrac{1}{\sec^2 x} = \cos^2 x \cdot \bf{Proved!} \)

mathslover (mathslover):

@elly394 , got it?

OpenStudy (anonymous):

Yes I did!! Thank you soooo much!!

mathslover (mathslover):

You're welcome :)